Theorem: Let X be normed liner space over a filed K and Let X' be its dual space of X. If X' is separable then X is separable
1
Expert's answer
2020-12-20T18:53:36-0500

Suppose that "X'" is separable. By definition, "\exists (f_n)_{n\in\mathbb{N}} \in (X')^{\mathbb{N}}, \overline{(f_n)} = X'" . For all "f_n\neq 0" we can find an "x_n \in X, f(x_n)\geq \frac{1}{2} ||f_n||" . We claim that "F = Vect_{\mathbb{Q}} ((x_n)_{n\in\mathbb{N}})" ("Vect_{\mathbb{Q}+i\mathbb{Q}}" if "K" is complex) is dense in "X". Suppose that it is not true, then "\exists x\in X, x\neq0, \epsilon>0, \forall y\in Vect_{\mathbb{Q}}((x_n)), ||x-y|| \geq \epsilon" . By Hahn-Banach theorem, "\exists f\in X', f(x) = 1 = ||f||, f(\overline{F})=f(F) = \{ 0\}" , as "f|_{\overline{Vect(x, (x_n))}}" is a continuous linear operator. Now let's consider "||f-f_n||" for "f_n\neq 0" :

"||f-f_n|| \geq \frac{|f(x_n)-f(x_n)|}{||x_n||} \geq \frac{1}{2}||f_n||"

But as "(f_n)" is dense, "\forall \epsilon'>0, \exists n" tel que "||f-f_n||<\epsilon' \Rightarrow ||f_n||<2\epsilon', ||f||<3\epsilon'". It is a contradiction, as "||f||=1."

For "f_n=0, ||f-f_n||=||f|| = 1 >0"

It is a contradiction, because "f \notin \overline{f_n}, f\in X'" . Therefore "F" is dense in "X" and thus "X" is separable.